I am bad at doing linear equations so i need help with a few homework peoblems. What is the equation of a line that goes through the point (4, 2) and is parallel to the line given by the equation y = 2x + 6?
Thanks for the help!

Found 2 solutions by macston, fractalier:
Answer by macston(5194)

Parallel lines have the same slope.
.
y=mx+b, m=slope
.
y=2x+6, m=slope=2
.
So parallel line will have the form: y=2x+b
.
To find b, use the point (4,2):
x=4 and y=2
.
y=2x+b
2=2(4)+b
2=8+b
-6=b Use this in the equation for b.
.
y=2x+b
y=2x-6
.
ANSWER: Equation for the parallel line through (4,2) is y=2x-6


Answer by fractalier(6550)

Okay, sure...
Parallel lines have equal slopes. That means the line you want has a slope of 2 also.
Now plug your info into y = mx + b and find b...here goes...
2 = (2)(4) + b
2 = 8 + b
b = -6
Now plug that into y = mx + b, with the slope you have and get
y = 2x - 6